OpenStudy (anonymous):

Please help!!! Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95 The answer choices are: A) 203 B) 197 C)168 D)163

OpenStudy (anonymous):

\[a _{1}=-7\] and \[a _{18}=95\]

OpenStudy (nincompoop):

do you know the formula for arithmetic sequence to get the nth term?

OpenStudy (anonymous):

No I do not :(

OpenStudy (nincompoop):

do you know how to use search engines?

OpenStudy (anonymous):

The formula is: \[ a_n = a_1+(n-1)d \]

OpenStudy (anonymous):

We can use \(a_1=-7\) and \(a_{18}=95\) to find \(d\).

OpenStudy (anonymous):

If we manipulate this equation, we find:\[ a_n=a_1+(n-1)d \implies a_n-a_1=(n-1)d \implies d=\frac{a_n-a_1}{n-1} \]

OpenStudy (jhannybean):

It was too much for her, she logged off.

OpenStudy (anonymous):

In general, we can say \[ d=\frac{a_n-a_k}{n-k} \]Where \(k<n\). The \(d\) corresponds to the slope.